OpenStudy (anonymous):

Write a 2 step equation that needs addition and multiplication in order to solve then solve it

OpenStudy (anonymous):

4x + 10 = 26

OpenStudy (anonymous):

there :)

OpenStudy (anonymous):

10 - 10 = 0 4x = 26 26 / 4 = 6.5 x = 6.5

OpenStudy (anonymous):

Kinda confusing in my opinion lol